# Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets

### v2

@@ Appendix D Additional experiments
Figure 8: Architectural design of AGFN. Top: The inductively biased parametrization of AGFN’s forward flow — based upon a GNN — enables the substantial reduction of the number of epochs required for training. Bottom: The use of a parametrized backward policy similarly enhances the training efficiency compared to a uniform policy. For both experiments, we considered $\mathcal{L}(\theta)<0.1$ as the early stopping criterion to interrupt AGFN’s training.
Trade-off between diversity and optimality in AGFN.
@@
as the reward upon which the GFlowNet is trained; if $T\rightarrow 0$, the distribution $p_{T}\propto R_{T}$ converges to a point mass in $R(\mathcal{G})$’s mode and, if $T\rightarrow\infty$, $p_{T}$ converges to an uniform distribution. This approach resembles the simulated tempering scheme commonly exploited in Monte Carlo methods Marinari and Parisi (1992) and was previously considered in the context of GFlowNets by Zhang et al. (2023). Figure 6 shows that progressively cold distributions (i.e., with $T\rightarrow 0$) lead to progressively concentrated and decreasingly diverse samples. Notably, the use of cold distributions may be adequate if we are highly confident in our score and are mostly interested in high-scoring samples (e.g., as in Rantanen et al., 2021).
Figure 9: Human-aided AGFN significantly outperforms alternative CD algorithms. Updating AGFN’s distribution according to the feedback of an oracle substantially improves AGFN’s capacity to correctly identify the true ancestral graph; indeed, a single feedback is sufficient to yield results better than (or indistinguishable from) alternative CD algorithms. We select the sampled AG with the highest posterior reward as a point estimate of AGFN and use the same datasets listed in table 2. The plots summarize the results of $30$ HITL simulations using $\pi=0.9$ and an oracle as an expert (see algorithm 1).
Sensitivity analysis for different noise levels.
@@
Figure 8 shows the increase in the training efficiency due to our architectural designs for parametrizing both the forward and backward flows of AGFN. Noticeably, the use of a two-layer graph isomorphism network (GIN; Xu et al., 2019) with a 256-dimensional embedding for the forward flow entailed a decrease of more than 10x in the number of epochs required for successfully training AGFN; this highlights the effectiveness of an inductively biased architectural design for the parametrization of GFlowNet’s flows. Correlatively, the use of a parametrized backward flow significantly enhances the training efficiency of AGFN and emphasizes the inadequacy of a uniformly distributed backward policy pointed out in a previous work Shen et al. (2023).
Human-aided AGFN versus alternative CD methods.
Figure 9 exposes the significant enhancement of AGFN’s point estimates entailed by our HITL framework for CD. This underlines the usefulness of the elicited knowledge, which is simply incorporated into our model through a re-weighting of the reward function, enabling the identification of the true ancestral graph. In contrast, most alternative CD algorithms cannot be as easily adapted to include various forms of expert knowledge — and such incorporation, when it is possible, usually precedes any inferential process Andrews (2020) or assumes the knowledge is perfect Wang et al. (2022).

### v3

@@ C.3 Sachs dataset details
Figure 10: Architectural design of AGFN. Top: The inductively biased parametrization of AGFN’s forward flow, based upon a GNN, enables the substantial reduction of the number of epochs required for training. Bottom: The use of a parametrized backward policy similarly enhances the training efficiency compared to a uniform policy. For both experiments, we considered $\mathcal{L}(\theta)<0.1$ as the early stopping criterion to interrupt AGFN’s training.
C.4 Learning a distribution over sparse AGs
To illustrate the efficiency and effectiveness of AGFN, we randomly select a 25-node graph and generate a 1000-sized dataset according to the corresponding SCM. Subsequently, we modified AGFN’s generative process by constraining the maximum degree a node can have to $3$, which drastically reduces the space of searchable graphs and is easily enforced by modifying the masking procedure described in Appendix B. Under these conditions, Figure 11 shows that AGFN accurately learns to sample from the target.
Figure 11: AGFN learns an accurate approximation to a distribution over 25-node sparse AGs. Left and right axes represent the same quantities at Figure 3’s left-most plot.
C.4.1 Trade-off between diversity and optimality in AGFN
@@
as the reward upon which the GFlowNet is trained; if $T\rightarrow 0$, the distribution $p_{T}\propto R_{T}$ converges to an uniform distribution over $R(\mathcal{G})$’s modes and, if $T\rightarrow\infty$, $p_{T}$ converges to an uniform distribution (Geman and Geman,, 1984, Theorem B). This approach resembles the simulated tempering scheme commonly exploited in Monte Carlo methods Marinari and Parisi, (1992) and was previously considered in the context of GFlowNets by Zhang et al., (2023). Figure 7 shows that progressively cold distributions (i.e., with $T\rightarrow 0$) lead to progressively concentrated and decreasingly diverse samples. Notably, the use of cold distributions may be adequate if we are highly confident in our score and are mostly interested in high-scoring samples (e.g., as in Rantanen et al.,, 2021).
C.5 Sensitivity analysis for different noise levels
@@ C.6 Ablation studies
Figure 10 shows the increase in the training efficiency due to our architectural designs for parametrizing both the forward and backward flows of AGFN. Noticeably, the use of a two-layer graph isomorphism network (GIN; Xu et al.,, 2019) with a 256-dimensional embedding for the forward flow entailed a decrease of more than 10x in the number of epochs required for successfully training AGFN; this highlights the effectiveness of an inductively biased architectural design for the parametrization of GFlowNet’s flows. Correlatively, the use of a parametrized backward flow significantly enhances the training efficiency of AGFN and emphasizes the inadequacy of a uniformly distributed backward policy pointed out in a previous work (Shen et al.,, 2023).
Yeast E. coli
N-ADMG $1762.50{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 448.51}} 22$$2217.61{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 228.28}}$
DCD $20941.02{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 127.62}}$ $21994.54{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 138.08}}$
GFCI $21000.43{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 183.23}}$ $21920.66{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 135.51}}$
FCI $1250.00{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 438.83}} 22$$2102.09{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 88.98}}$
ACI $1215.77{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 246.04}} 22$$2071.27{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 188.83}}$
AGFN $\mathbf{20918.87}{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 71.89}}$ $\mathbf{21852.23}{\scriptstyle{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@gray@stroke{0.5}\pgfsys@color@gray@fill{0.5}\pm 45.29}}$
Table 3: BIC for GRN datasets. AGFN’s AGs have significantly larger scores (smaller BIC) than baselines’.
C.7 Experiments on gene regulatory networks (GRN)
Table 3 shows that AGFN leads to significantly better results in terms of BIC than some of the considered baselines in the structure learning of the 10-node GRNs of a yeast of the E. coli bacteria. This observation underlines the effectiveness of our method in real-world datasets. Moreover, the results in Table 3 attest that AGFN may be successfully scaled beyond the relatively small 5-node causal diagrams considered in the main body of text. For this experiment, we implemented a GFlowNet using the same architecture outlined in section C.2 to sample proportionally to the BIC score based on 200 data points. The datasets were obtained from the DREAM3 challenge.
C.8 Assessment of our elicitation framework
Figure 12 shows that our entropy-minimizing elicitation strategy leads to consistently faster reduction in the BIC when compared against an algorithm that selects the next query uniformly at random. In practice, this emphasizes that our approach for probing the expert significantly improves the convergence speed of the expert-in-the-loop process towards better (higher-scoring) AGs. In addition, the algorithm we propose incurs a minimal bottleneck to the iterative refinement of a trained AGFN. That said, the development of provably optimal experimental designs for probing an expert in the context of CD remains open.
chain4 IV collfork
BIC
[…]
Table 4: Average BIC. The (above the dashed line) denotes methods yielding point estimates, for which we employ bootstrap to report the mean and average standard deviation. For N-ADMG and AGFN, we estimate the quantities using 100100k sampled graphs. In most cases, AGFN incurs higher sample diversity than bootstrapped non-probabilistic CD, and achieves comparable results to these methods in terms of BIC (lower is better).
Figure 12: CD with simulated human feedback on synthetic graphs. Top/bottom row shows the mean SHD/BIC of AGFN samples as a function of human interactions. Probing the expert about the edge that minimizes the mean cross-entropy leads to a faster decrease in BIC than a random strategy. SHD decreases similarly in both cases. Results reflect the outcomes of 30 simulations.
Appendix D Details on the Experiments with LLMs
For the preliminary experiments, we used both GPT-4o (OpenAI et al.,, 2024) and Deepseek (V3 and R1) (DeepSeek-AI et al.,, 2024). However, Deepseek’s outputs were too far from the correct edges. As the feedback was misleading the algorithm, the final results were worse than using AGFN without any feedback. As GPT-4o’s feedback was much closer to the real edges, though far from perfect, we decided to report only these results.
Our LLM-based framework is built using DSPy (Khattab et al.,, 2022, 2024), a library to build modular AI software. We define to the LLM what the input and output will be:
description=(
"Pair of nodes about which we want to know the relationship (defined by EdgeType). "
"Each pair of nodes may be directly associated, via direct causation, or indirectly associated, via latent confounding. "
"The nature of each node and the underlying problem is described in the context field."
context: str = dspy.InputField(
description=(
"The context of the problem, including the nature of each node (represented by a string) "
"and the nature of the kind of relationships we are looking for."
relationship: EdgeType = dspy.OutputField(
description=(
"This is the feedback. It describes the nature of the relationship between the input nodes. "
"This information should be inferred from the provided context, background knowledge, and the nature of the problem."
Then, for each given query, we prompt the LLM:
‘You are an expert on the human immune system cell. You are investigating the cause-and-effect relationships between a specific set of observed variables representing proteins and phospholipids: "Raf", "Mek", "Plcg", "PIP2", "PIP3", "Erk", "Akt", "PKA", "PKC", "P38", "Jnk". Your task is to determine the causal relationship between the given variables. If there are both indirect and direct relationships between the variables, you should describe only the direct one.’
We repeat the process $10$ times and use the most answered feedback. As its uncertainty, we consider the simple ratio of the number of times the LLM gave that feedback over $10$.
D.1 Elicitation Framework
When considering an LLM as an expert, we extend our elicitation framework to accommodate a query-dependent confidence level. To understand this, recall that our refinement algorithm iteratively collects a feedback $f_{r}\in\{1,2,3,4\}$ regarding a selected edge $r$ from an expert with confidence $\pi$. In particular, our model in Section 4 assumes that $\pi$ does not depend on $r$. To relax this constraint, we allow for a $r$-dependent likelihood function in the hierarchical Bayesian model defining our expert-in-the-loop process. That is, our model becomes
$\omega_{r}$ $\sim\text{Cat}(\boldsymbol{\rho}_{r}),$ (29)
$f_{r}|\omega_{r}$ $\sim\text{Cat}\left(\delta_{\omega_{r}}\cdot\pi_{r}+(\mathbf{1}-\delta_{\omega_{r}})\cdot\left(\frac{1-\pi_{r}}{3}\right)\right)$ (30)
(using the same notation as in Section 4). Although demonstrably effective (see our experiments in Section 5), this approach deviates our model from a purely Bayesian paradigm. We leave the investigation of a fully Bayesian approach for knowledge elicitation in causal discovery to future endeavors. For example, this could be achieved through the introduction of a Beta prior distribution over $\pi$.
Also, as explained in the previous section, we elicit $\pi_{r}$ from the LLM through a bootstrapping algorithm. In doing so, we draw on recent studies on the calibration of the verbalized confidence of LLMs (Zhang et al.,, 2024; Xiong et al.,, 2024; Yang et al.,, 2024; Tao et al.,, 2024). Xiong et al., (2024), for example, shows that LLMs often overstate their confidence and that averaging the LLM’s responses across multiple queries can provide more accurate uncertainty estimates. Interestingly, our early (unreported) experiments using the LLM’s verbalized confidence were also less successful than the ones that relied on bootstrapping.
Appendix E Related Work
E.1 CD under latent confounding.
Following the seminal works by Spirtes et al., (2001) and Zhang, 2008b introducing the complete FCI, a variety of works have emerged. Among them are algorithms designed for sparse scenarios, including RFCI (Colombo et al.,, 2012) and others (Silva,, 2013; Claassen et al.,, 2013). Notably, Silva, (2013)’s framework uses a Bayesian approach to CD of Gaussian causal diagrams based on sparse covariance matrices. Nonetheless, it requires sampling one edge at a time and relies on numerical heuristics that might effectively alter the posterior we are sampling from. Colombo et al., (2012) introduced the conservative FCI to handle conflicts arising from statistical errors in scenarios with limited data, even though it yields less informative results. Subsequent efforts to improve reliability led to the emergence of constraint-based CD algorithms based in Boolean satisfiability (Hyttinen et al.,, 2014; Magliacane et al.,, 2016), although they are known to scale poorly on $|\mathbf{V}|$ (Lu et al.,, 2021). In another paradigm, score-based search algorithms rank MAGs according to goodness-of-fit measures, commonly using BIC for linear Gaussian SCMs (Triantafillou and Tsamardinos,, 2016; Zhalama et al., 2017a, ; Rantanen et al.,, 2021). There are also hybrid approaches that combine constraint-based strategies to reduce the search space, such as GFCI (Ogarrio et al.,, 2016), M3HC (Tsirlis et al.,, 2018), BCCD (Claassen and Heskes,, 2012) and GSPo (Bernstein et al.,, 2020). Continuous optimization has recently emerged as a novel approach to score-based CD, as DCD (Bhattacharya et al.,, 2021) and N-ADMG (Ashman et al.,, 2023). While N-ADMG focuses on a more restricted setting compared to AGs, it offers some uncertainty quantification in the variational posterior, making it the most directly comparable baseline. We rigorously follow the protocols outlined in the original works.
E.2 CD with expert knowledge.
Previous works on CD have explored various forms of background knowledge. This includes knowledge on edge existence/non-existence (Meek,, 1995), ancestral constraints (Ribeiro et al.,, 2024; Chen et al.,, 2016), variable grouping (Parviainen and Kaski,, 2017), partial order (Andrews,, 2020) and typing of variables (Brouillard et al.,, 2022). Incorporating expert knowledge is pivotal to reducing the search space and the size of the learned equivalence class. Nevertheless, due to significant challenges, up to date, there are only a few works trying to integrate background knowledge into CD within the context of latent confounding (Andrews,, 2020; Wang et al.,, 2022). These works operate under the assumption of perfect expert feedback. In contrast, our contribution is the first to deal with more realistic situations where expert input might be inaccurate.
Recently, many works have been using the domain knowledge within LLMs to improve CD algorithms. Usually, the frameworks focus on using the LLM alongside a data-driven model. To achieve this, specialized querying strategies are often employed (Wan et al.,, 2024), for example,
  1. 1.
    Pair-wise discovery;
  2. 2.
    Conditional independence tests;
  3. 3.
    Full graph discovery.
Most works elicit knowledge from the LLM in pairs of variables (Ban et al.,, 2025; Jiralerspong et al.,, 2024). However, there are some that focus on discovering the full graph at once (Ban et al.,, 2023). On top of that, recent works have also used LLMs for providing indirect information that improves the CD process. For example, i) Li et al., (2025) uses the LLM for intervention targeting, while ii) Vashishtha et al., (2023) solely focus on finding a causal order, instead of the causal graph. Remarkably, none of the mentioned works considers any uncertainty metric for the answers given by the LLM, which we do.
A recent paper by Ankan and Textor, (2025) also considered an expert-in-the-loop which could be either a human or an LLM. The approach is based on using CI statements or an iteration of the classical PC algorithm (Spirtes et al.,, 1993) together with an expert-in-the-loop. Differently from our work, this outputs a single DAG and does not consider any uncertainty over the final result. Moreover, there are experiments using LLMs as experts. Importantly, the used dataset does not have a defined true graph to be compared to and the code is not available. So, we were unable to test their framework and use it as comparison.
E.3 Generative Flow Networks
GFlowNets (Bengio et al., 2021a, ) are generative models that sample discrete composite objects from an unnormalized reward function. They have been successfully used to a variety of structures, including protein sequences (Jain et al.,, 2022), molecules (Bengio et al., 2021a, ), schedules (Zhang et al.,, 2023), phylogenetic trees (Anonymous,, 2023), strategies in adversarial games (Jiralerspong et al.,, 2023), dropout masks (Liu and et al.,, 2023), sentences (Hu et al., 2023a, ). They have also been used to train energy-based models (Zhang et al.,, 2022; Hu et al., 2023b, ). Recently, Garipov et al., (2023) presented a strategy for reusing expensively pretrained GFlowNets based on composition and classifier guidance. In the field of structure learning, they have been applied to Bayesian networks, more specifically to sample a posterior over DAGs, although without accounting for unobserved confounding (Deleu et al.,, 2022), and to dynamic Bayesian networks (Atanackovic et al.,, 2023). Recently, Deleu et al., (2023) proposed an extension to jointly infer the structure and parameters, also grounded in the assumption of causal sufficiency. It is worth highlighting that training GFlowNets in these scenarios presents optimization challenges, resulting in the utilization of a variety of loss functions (Shen et al.,, 2023; Malkin et al.,, 2022; Pan et al., 2023a, ; Pan et al., 2023b, ). Moreover, Lahlou et al., (2023) extended GFlowNets to continuous domains.
Appendix F Details on (Non-)Ancestral Knowledge
Our EITL approach builds on the premise that experts can determine whether a variable is a cause (ancestor) of another, regardless of whether this causation is direct or confounded. This knowledge is much coarser compared to that needed for a causal DAG. While directed edges in a causal DAG represent direct causation (assumed to be unconfounded), which can be challenging for experts to identify, edges in an AG represent only the existence of a directed (causal) path, regardless of any other potentially unobserved, confounding paths.
Scientists often have an understanding of the ancestral (or non-ancestral) relationship between two variables. Illustratively, in biomedicine, causal knowledge primarily arises from randomized experiments, readily available on platforms like PubMed (Wheeler et al.,, 2007), DrugBank (Knox et al.,, 2024), and GO-CAM (Thomas et al.,, 2019). Yet, it is important to note that causality from experimental studies only guarantees ancestrality, not directness or unconfoundedness. For example, causal effects may be mediated by unobserved factors, and confounding, albeit reduced by randomization in experiments, can persist in observational studies. An expert can incorporate these established ancestral relations using our framework by increasing the probability of the corresponding directed edges.
There are also numerous examples of known non-ancestral relationships. For instance, it is widely accepted that sociodemographic variables (e.g., age and sex), as well as genetic variables (referred to as $X$), are not caused by factors such as drugs, diseases, or other phenotypes (referred to as $Y$). To reflect this understanding that $Y$ is not an ancestor of $X$, one can reduce the likelihood of $Y$ being ancestral to $X$, achieved by decreasing the probability of the corresponding edges $Y\rightarrow X$.
Additionally, information indicating that an association is purely spurious, with no causation in any direction, may also be accessible. For instance, the link between coffee consumption ($X$) and heart disease ($Y$) is likely to be solely spurious, as factors such as smoking, diet, and lifestyle habits among heavy coffee drinkers could confound the results. This knowledge is incorporated by increasing the probability of a bidirected edge $X\leftrightarrow Y$.
Finally, our expert-in-the-loop approach considers the confidence level of such knowledge, facilitating its inclusion even when the source may lack reliability.